how do I solve this equation?
-0.3(x-6)+ x = 0.5(3-x)

Answer by rfer(16322)

-0.3x+1.8+x=1.5-0.5x
-0.3x+0.5x+x=1.5-1.8
1.2x=-0.3
x=-0.3/1.2
x=-0.25